OpenStudy (anonymous):

Write an equation of a circle with the given center and radius. Center (4,5), radius 7. A. (x-4)^2+(7-5)^2=7 B. (x-4)^2+y-5)^2=49 C. (x+4)^2+(y-5)^2=7 D. (x+4)^2+(y-5)^2=49 I will medal

OpenStudy (luigi0210):

You're going to need to use \(\Large (x-h)^2+(y-k)^2=r^2 \) Where \(\large (h, k) \) is the center and \(\large r\) is the radius

OpenStudy (anonymous):

so (x-4)^2+(y-5)^2=7^2= 49?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Yup :)